OpenStudy (anonymous):

how can i find the polynomial function with roots 1, 7, and –3 of multiplicity 2 ? my notes dont explain this very well

OpenStudy (ash2326):

We are given that the polynomial has roots 1, 7 and -3. Each with multiplicity of 2 so the polynomial can be written as \[(x-1)^2(x-7)^2(x+3)^2=0\] Here x-1 is a factor , root 1 Has a multiplicity of 2 so (x-1)^2 , similarly for other roots. :)

OpenStudy (ash2326):

Did you get wow?

hero (hero):

I give you the medal under the assumption that multiplicity 2 is correct.

OpenStudy (ash2326):

It's given in the question.

OpenStudy (anonymous):

\[(x+1)(x+49)(x+9)\]

OpenStudy (anonymous):

would i end up with that?

OpenStudy (ash2326):

Use this formula \[(x-a)^2=x^2-2ax+a^2)\]

OpenStudy (ash2326):

\[(x-1)^2=x^2-2x+1\] do this for all the other terms

OpenStudy (anonymous):

i got a 3

OpenStudy (ash2326):

How did you get 3?

OpenStudy (anonymous):

\[(x^{2}-2(-1)(1) +-1 ^{2})\]

OpenStudy (anonymous):

x^2+2+1=3

OpenStudy (ash2326):

wow it'll be \[\huge{ x^2-2(1)(x)+1=x^2-2x+1}\]

OpenStudy (ash2326):

\[\huge{(x-2)^2=x^2-2(x)(2)+2^2}\]

OpenStudy (ash2326):

Do this for the question, post your answer. I'll help if you are stuck somewhere

OpenStudy (anonymous):

is it 4?

OpenStudy (ash2326):

nope it's \[x^2-2(x)(2)+2^2=x^2-4x+4\]

OpenStudy (anonymous):

so then what do i do next?

OpenStudy (ash2326):

we have \[(x-1)^2(x-7)^2(x+3)^2=0\] expand each bracketed term, and then leave, no need for multiplying them, It'll be very long \[(x^2-2x-1)(x-7)^2(x+3)^2=0\] expand (x-7)^2 and (x+3)^2 also and that'll be your answer

OpenStudy (anonymous):

(x^2 - 14x + 49)(x^2 + 6x + 9) ?

OpenStudy (ash2326):

yeah correct:)

OpenStudy (anonymous):

thank you

OpenStudy (ash2326):

welcome :)